# HarmonRank: Ranking-aligned Multi-objective Ensemble for Live-streaming E-commerce Recommendation

### v1

@@ 1. Introduction
To balance these competing objectives, industrial ranking systems usually follow a two-stage paradigm (see Fig. 1 (a)), i.e., multi-objective prediction (MP) and multi-objective ensemble (ME) (ali_ltr; pantheon; intel; label_aggregation). The MP stage is usually a high-capacity multi-task learning model (mmoe) whose goal is to predict the precise probability scores for multiple objectives, e.g., click through rate (CTR). Whereas ME stage is a light-weight model whose goal is to generate a unified ensemble score in a personalized manner. It usually only takes aforementioned multi-objective scores and a few user features as inputs. The ensemble score is used to truncate top items to represent for users. In this paper, we focus on ME problem. Similar to MP stage, in ME stage we use AUC (Area Under ROC Curve) (auc) between the ensemble score and the ground-truth labels to evaluate the ranking abilities of the ensemble model. The only difference is that we utilize the sum of AUC of all objectives as the metic for comprehensive multi-objective ranking evaluation.
The key challenge at this stage lies in the absence of a gold-standard supervision to comprehensively assess items considering multiple objectives. To mitigate this problem, a typical solution is to impose multiple independent binary classification supervisions over the ensemble score (pantheon; peltr) (see Fig. 1 (b)). However, this paradigm suffers from two limitations, for which we present a conceptual view in Fig. 1 (c). (1) Lack of alignment to ranking task. The optimization directions of binary classification based supervisions are misaligned with the ultimate ranking task (evaluated by AUC) (tang2022smooth). For an example, if the positive labels of some action are highly sparse, a dummy predictor which predicts all instances as negative seems to present an acceptable binary classification result. But it actually yields a poor ranking results with an AUC of 0.5. (2) Lack of inter-objective alignment. The different objectives are neither independent nor strictly adversarial - they are partially dependent. For an example, buy actions empirically exhibit high correlations with comment actions, which can be verified through the statistical analysis of correlation between objectives (see Fig.4). Thus, if model can learn the aligned parts among multi-objectives, it is easier for model to trade-off these competing objectives.
@@ 2.1. Multi-objective optimization in recommender systems
Multi-objective optimization is recognized as a key challenge in both research and industry community. Early recommender systems primarily optimized for only few objectives, e.g. click-through rate (CTR) prediction for e-commerce recommender system and watch time for video recommender systems. However, with the increasing complexity user interface functionalities, morden recommender system need to balance more competing objectives. Some studies attend on the sample selection bias caused by chain relationships between objectives, e.g., CTR and conversion rate (CVR) (essm; choruscvr). Some other researchers are dedicated to improve multi-task learning mechanism between different objectives (mmoe; ple). For an example, Multi-gate Mixture-of-Experts (MMOE) (mmoe) architecture addresses this by leveraging expert networks to capture shared representations across objectives, thereby improving accuracy of individual objectives. Despite these advances, how to ensemble multiple prediction scores of different objectives remains a non-trivial challenge. Current industrial systems typically address this through a multi-objective ensemble model (intel; unsupervised; ali_ltr).
@@ 2.2. Multi-objective Ensemble in recommender systems
Early industrial approaches predominantly employed ranking formulas to combine scores from multiple objectives, typically optimizing parameters through grid search. While black-box optimization methods, e.g., Bayesian optimization (bayesian) and cross entropy methods (cross_entropy), can reduce the iterations for searching optimum, they still suffer from two critical limitations: (1) inability to present personalized ranking mechanism to adapt user preferences, and (2) incapacity to exploit real-time user feedback. The factors matter for system since the optimal ensemble schemes change with different users and different time. To address these challenges, recent methods have proposed model-based ensemble solutions, to utilize more informative features and deep neural networks trained in hourly updated streaming feedback.
The key challenge for model-based multi-objective ensemble lies in the absence of a gold-standard supervision across multiple objectives. To solve this problem, some methods apply multiple unilateral binary classification losses to optimize towards all objectives simultaneously (pantheon; peltr), which can be referred as loss aggregation methods. The other methods aim to build a unified supervision by aggregating multiple binary labels into a single regression label (intel; label_aggregation), which can be categorized as label aggregation methods. Despite these efforts, the optimization directions of these methods’ supervisions are misaligned with the ultimate ranking task (evaluated by AUC), which leads to suboptimal performance. Reinforce learning based methods (ali_ltr) can directly optimize towards ranking task metrics (AUC sum), whereas their policy gradients suffer from large variance and instability. Few of existing methods consider to optimize final sum of AUC in an end-to-end differentiable manner. It is NP-hard to direct optimize AUC because of the non-convexity and discontinuousness of (auc_pl).
@@ 2.3. AUC optimization
The AUC is a ranking quality metric defined based on pairwise comparisons (auc; hanley1982meaning) under binary label settings. Thus, traditional differentiable AUC optimization methods concentrate on developing effective pairwise surrogate functions to approximate non-differentiable comparison function, i.e., $s_{i}-s_{j}\geq 0$. For examples, pairwise square methods (psq) employ square functions $(1-(s_{i}-s_{j}))^{2}$ surrogate functions, while pairwise logistic methods (auc_pl) propose to use logistic function $\text{log}(1+\text{exp}(-(s_{i}-s_{j})))$ to approximate the comparison function. However, the inherent $O(n^{2})$ time complexity of pairwise approaches renders them impractical for large-scale datasets. Consequently, numerous optimization attempts have shifted towards instance-wise AUC optimization. Leveraging the special form of pairwise square, pairwise AUC maximization can be decomposed into a min-max instance-wise formulation (ying2016stochastic), with $O(N)$ time complexity. Nevertheless, this instance-wise optimization introduces much sensitivity to noisy samples. To address this, AUCM (aucm) proposed a margin-based min-max optimization strategy to effectively mitigates this issue.
@@ 4.1.1. Rank-based AUC formulation
Since $\mathcal{G}$ is non-differentiable w.r.t. $\mathbf{\theta}$, existing wisdom usually maximize AUC via surrogate function to approximate pairwise comparing function $\mathbb{I}(s_{i}\geq s_{j})$ i.e., pairwise logistic (auc_pl) and pairwise square (psq) functions. However, this formulation brings large inconsistency between physical AUC computation. Meanwhile, pairwise computation introduce quadratic computation cost during training, which hinders further scaling on datasets. To avoid these problems, we resort to alternative formulation for AUC computation.
In this section, we temporarily omit the subscripts $m$ of objectives for simplicity. When sorting all positive and negative samples in an increasing order, AUC can be simplified to compute the rank sum of positive samples for the good property of ordered list (rank_sum_test_v2):
@@
However, for rank operation $\text{Rank}(\cdot)$ is non-convex, discontinuous gradients is infeasible for $\mathbf{s}$. To alleviate this, we employ a fast differentiable ranking algorithm (fast_soft_sort) to enable gradient backpropagation from supervisions to the ensemble model.
@@ 4.1.2. Differentiable Ranking
Following (fast_soft_sort), we cast ranking operation (in increasing order here) as a discrete optimization problem over all feasible permutations $\Sigma$.
@@
In this way, the original problem has been transformed to the projections to the permutahedron, which is a strong convex function. This approach enables forward propagation with $O(n\text{log}n)$ time complexity for sorting operation. In the same time, it cost only $O(n)$ time complexity during backward propagation for the gradients computation can be finished in the sorted sequence (fast_soft_sort).
@@ 4.2.1. Relation-aware Module
We align the shared parts between objectives with a relation-aware two-step paradigm, first align then ensemble. For align step, as shown in Fig. 2(b), we utilize self-attention mechanism (self_attention) to capture pairwise relations between objectives. Through dynamic computation of relations between objectives, the aligned common parts between them are extracted. Specifically, the softmax normalized dot products between projected objective vectors, $\mathbf{Q}_{r}$ and $\mathbf{K}_{r}$, form the self attention matrix $\mathbf{A}_{r}\in\mathbb{R}^{m\times m}$ between them. This matrix naturally represents the relation strengths between objectives. Then the original score embeddings are integrated into the representations $\mathbf{x}_{r}$ according to the attention matrix.
@@ 4.3. Pre-processing
Before feed the score into relation-aware and relation-agnostic modules. We apply discretization embedding to represent the multi-objective scores to obtain non-linear representation. We use a simple equal distance discretization embedding technique to transform scalar scores to embeddings (guo2021embedding).
@@ 5. Experiments
Multi-objective BCE (pantheon) 2.2420 3.6465 2.1555 3.6682
Label aggregation (intel) 2.4040 3.8311 2.0765 3.6883
Reinforce learning (ali_ltr) 2.2488 3.9074 2.1930 3.7467
@@ 5.1. Implementation Details
Before training ensemble model, we firstly prepare the input multi-objective scores via performing inference with a high-capacity ranking model over both datasets. For Kuaishou-Elive, we use the online in-service model trained incrementally by years of logs. For the public TenRec dataset, we train a standard MMOE (mmoe) ranking model. Based on prepared multi-objective scores, all offline models are trained with 500 epochs with SGD optimizer. In terms of online training model with streaming data, it is usually trained for one epoch for the consideration of run-time cost and one-epoch overfiting risk in ranking stage. This scheme results in low sample efficiency. Since the model of ME stage is much more lightweight than MP stage and no sparse ID feature is used (cause of one-epoch overfitting), we train 20 epochs for each batch of streaming samples, to balance run-time cost and sample efficiency. We set learning rate to 1e-4 and batch size = 10240 for all methods.
@@ 5.2. Datasets
Since there are few public live-streaming e-commerce dataset, we build a new dataset called Kuaishou E-Live, which contains both e-commerce feedback (purchase) and user engagement oriented feedbacks (long view, like, follow and comment). We built it by sampling millions of exposure logs at Kuaishou e-commerce platform, which is a pioneering live-streaming e-commerce platform with over 400 million DAUs. We utilize several context and user profile features including the hour of a day, app version, age and gender, etc., to characterize user preference in ensemble learning.
Moreover, to verify the universality of over method, we conduct experiments on a public short video recommendation dataset, TecRec-QKVideo (yuan2022tenrec). TecRec-QKVideo is a collected from QQ-KAN video feeds platforms. It contains over 1M user interaction logs including five type of feedbacks, click, like, share, follow and long view. For the need of personalized ensemble, we utilize several user profile features like age, gender, etc. We train our method and compared methods with the same equal loss weights (all weights set to 1.0) for fair comparison on objective trade-offs. We refer more details about dataset to appendix section A.
@@
  • Multi-objective binary cross entropy based (M-BCE) (pantheon): This method enables optimization towards multiple objectives by integrate joint learning under multiple independent binary cross entropy losses for different objectives.
  • Reinforce learning (RL) based (ali_ltr) : RL-based method formulate multi-objective ensemble as a markov decision process, and treat input multi-objective scores as the ‘state’ and the multi-object fusion score as the ‘action’. The multi-object AUC over a batch of data serve as ‘reward’.
  • Label aggregation based (label_aggregation; intel) : Label aggregation based method transforms multi-objective learning into single-objective learning by fusion multiple binary labels to a unified regression label. Then it utilize mean square error loss between the ensemble score and the aggregated label to optimize the ensemble model.
@@ 5.4. Offline Results (RQ1)
The experimental results on two datasets are shown in Tab. 1. On the TenRec-QKVideo dataset, our proposed method outperforms MBCE (pantheon) by 27.2 pp (3.919 v.s. 3.647) for 5-objective AUC sum. On the Kuaishou-Elive dataset, our model outperforms MBCE by 8.4 pp (3.7522 v.s. 3.6682) for 5-objective AUC sum. The comparison results to the MBCE model demonstrates the efficacy of our proposed ranking-aligned ensemble model.
The RL-based model is the best performing baseline on both datasets. On the TenRec-QKVideo dataset, our model outperforms RL-based model (ali_ltr) by 1.2 pp (3.919 v.s. 3.907) for 5-objective AUC sum. On the Kuaishou-ELive dataset, our model outperforms RL-based model by 0.55 pp (3.7522 v.s. 3.7467). These results verify our continuous gradient-based method is superior to RL policy gradient based methods. In a nutshell, our model consistently outperforms the best-performing baselines on both two datasets in a large margin in terms of different experimental settings.
@@ 5.5. Ablation study (RQ2)
Other than our proposed differentiable ranking based AUC maximization methed, we compare several popular pairwise AUC surrogate losses i.e., pairwise logistic (auc_pl), pairwise square (psq) and the state-of-the-art instance-wise AUC maximization method AUCM (aucm). Pairwise methods maximize AUC objective with logistic or square function surrogate to approximate pairwise $\mathbb{I}(f(\mathbf{x}_{+})\geq f(\mathbf{x}_{-}))$. For pairwise computation nature, these methods perform a $O(n^{2})$ time complexity during training process. Different from them, AUCM transforms pairwise maximization objective to an equivalent instance-wise min-max problem, which reduce the training time complexity to $O(n)$. Notwithstanding these efforts, the inherent inconsistency with the AUC metric persists.

### v2

@@ 1. Introduction
To balance these competing objectives, industrial ranking systems usually follow a two-stage paradigm (see Fig. 1 (a)), i.e., multi-objective prediction (MP) and multi-objective ensemble (ME) (Pei et al., 2019; Cao et al., 2025; Li et al., 2023; Carmel et al., 2020). The MP stage is usually a high-capacity multi-task learning model (Ma et al., 2018a) whose goal is to predict the precise probability scores for multiple objectives, e.g., click through rate (CTR). Whereas ME stage is a light-weight model whose goal is to generate a unified ensemble score in a personalized manner. It usually only takes aforementioned multi-objective scores and a few user features as inputs. The ensemble score is used to truncate top items to represent for users. In this paper, we focus on ME problem. Similar to MP stage, in ME stage we use AUC (Area Under ROC Curve) (Bamber, 1975) between the ensemble score and the ground-truth labels to evaluate the ranking abilities of the ensemble model. The only difference is that we utilize the sum of AUC of all objectives as the metic for comprehensive multi-objective ranking evaluation.
The key challenge at this stage lies in the absence of a gold-standard supervision to comprehensively assess items considering multiple objectives. To mitigate this problem, a typical solution is to impose multiple independent binary classification supervisions over the ensemble score (Cao et al., 2025; Lin et al., 2019) (see Fig. 1 (b)). However, this paradigm suffers from two limitations, for which we present a conceptual view in Fig. 1 (c). (1) Lack of alignment to ranking task. The optimization directions of binary classification based supervisions are misaligned with the ultimate ranking task (evaluated by AUC) (Tang et al., 2022). For an example, if the positive labels of some action are highly sparse, a dummy predictor which predicts all instances as negative seems to present an acceptable binary classification result. But it actually yields a poor ranking results with an AUC of 0.5. (2) Lack of inter-objective alignment. The different objectives are neither independent nor strictly adversarial - they are partially dependent. For an example, buy actions empirically exhibit high correlations with comment actions, which can be verified through the statistical analysis of correlation between objectives (see Fig.4). Thus, if model can learn the aligned parts among multi-objectives, it is easier for model to trade-off these competing objectives.
@@ 2.1. Multi-objective optimization in recommender systems
Multi-objective optimization is recognized as a key challenge in both research and industry community. Early recommender systems primarily optimized for only few objectives, e.g. click-through rate (CTR) prediction for e-commerce recommender system and watch time for video recommender systems. However, with the increasing complexity user interface functionalities, morden recommender system need to balance more competing objectives. Some studies attend on the sample selection bias caused by chain relationships between objectives, e.g., CTR and conversion rate (CVR) (Ma et al., 2018b; Cheng et al., 2025). Some other researchers are dedicated to improve multi-task learning mechanism between different objectives (Ma et al., 2018a; Tang et al., 2020). For an example, Multi-gate Mixture-of-Experts (MMOE) (Ma et al., 2018a) architecture addresses this by leveraging expert networks to capture shared representations across objectives, thereby improving accuracy of individual objectives. Despite these advances, how to ensemble multiple prediction scores of different objectives remains a non-trivial challenge. Current industrial systems typically address this through a multi-objective ensemble model (Li et al., 2023; Yu et al., 2024; Pei et al., 2019).
@@ 2.2. Multi-objective Ensemble in recommender systems
Early industrial approaches predominantly employed ranking formulas to combine scores from multiple objectives, typically optimizing parameters through grid search. While black-box optimization methods, e.g., Bayesian optimization (Močkus, 1974) and cross entropy methods (Rubinstein and Kroese, 2004), can reduce the iterations for searching optimum, they still suffer from two critical limitations: (1) inability to present personalized ranking mechanism to adapt user preferences, and (2) incapacity to exploit real-time user feedback. The factors matter for system since the optimal ensemble schemes change with different users and different time. To address these challenges, recent methods have proposed model-based ensemble solutions, to utilize more informative features and deep neural networks trained in hourly updated streaming feedback.
The key challenge for model-based multi-objective ensemble lies in the absence of a gold-standard supervision across multiple objectives. To solve this problem, some methods apply multiple unilateral binary classification losses to optimize towards all objectives simultaneously (Cao et al., 2025; Lin et al., 2019), which can be referred as loss aggregation methods. The other methods aim to build a unified supervision by aggregating multiple binary labels into a single regression label (Li et al., 2023; Carmel et al., 2020), which can be categorized as label aggregation methods. Despite these efforts, the optimization directions of these methods’ supervisions are misaligned with the ultimate ranking task (evaluated by AUC), which leads to suboptimal performance. Reinforce learning based methods (Pei et al., 2019) can directly optimize towards ranking task metrics (AUC sum), whereas their policy gradients suffer from large variance and instability. Few of existing methods consider to optimize final sum of AUC in an end-to-end differentiable manner. It is NP-hard to direct optimize AUC because of the non-convexity and discontinuousness of (Gao and Zhou, 2012).
@@ 2.3. AUC optimization
The AUC is a ranking quality metric defined based on pairwise comparisons (Bamber, 1975; Hanley and McNeil, 1982a) under binary label settings. Thus, traditional differentiable AUC optimization methods concentrate on developing effective pairwise surrogate functions to approximate non-differentiable comparison function, i.e., $s_{i}-s_{j}\geq 0$. For examples, pairwise square methods (Gao et al., 2013) employ square functions $(1-(s_{i}-s_{j}))^{2}$ surrogate functions, while pairwise logistic methods (Gao and Zhou, 2012) propose to use logistic function $\text{log}(1+\text{exp}(-(s_{i}-s_{j})))$ to approximate the comparison function. However, the inherent $O(n^{2})$ time complexity of pairwise approaches renders them impractical for large-scale datasets. Consequently, numerous optimization attempts have shifted towards instance-wise AUC optimization. Leveraging the special form of pairwise square, pairwise AUC maximization can be decomposed into a min-max instance-wise formulation (Ying et al., 2016), with $O(N)$ time complexity. Nevertheless, this instance-wise optimization introduces much sensitivity to noisy samples. To address this, AUCM (Yuan et al., 2021) proposed a margin-based min-max optimization strategy to effectively mitigates this issue.
@@ 4.1.1. Rank-based AUC formulation
Since $\mathcal{G}$ is non-differentiable w.r.t. $\mathbf{\theta}$, existing wisdom usually maximize AUC via surrogate function to approximate pairwise comparing function $\mathbb{I}(s_{i}\geq s_{j})$ i.e., pairwise logistic (Gao and Zhou, 2012) and pairwise square (Gao et al., 2013) functions. However, this formulation brings large inconsistency between physical AUC computation. Meanwhile, pairwise computation introduce quadratic computation cost during training, which hinders further scaling on datasets. To avoid these problems, we resort to alternative formulation for AUC computation.
In this section, we temporarily omit the subscripts $m$ of objectives for simplicity. When sorting all positive and negative samples in an increasing order, AUC can be simplified to compute the rank sum of positive samples for the good property of ordered list (Hanley and McNeil, 1982b):
@@
However, for rank operation $\text{Rank}(\cdot)$ is non-convex, discontinuous gradients is infeasible for $\mathbf{s}$. To alleviate this, we employ a fast differentiable ranking algorithm (Blondel et al., 2020) to enable gradient backpropagation from supervisions to the ensemble model.
@@ 4.1.2. Differentiable Ranking
Following (Blondel et al., 2020), we cast ranking operation (in increasing order here) as a discrete optimization problem over all feasible permutations $\Sigma$.
@@
In this way, the original problem has been transformed to the projections to the permutahedron, which is a strong convex function. This approach enables forward propagation with $O(n\text{log}n)$ time complexity for sorting operation. In the same time, it cost only $O(n)$ time complexity during backward propagation for the gradients computation can be finished in the sorted sequence (Blondel et al., 2020).
@@ 4.2.1. Relation-aware Module
We align the shared parts between objectives with a relation-aware two-step paradigm, first align then ensemble. For align step, as shown in Fig. 2(b), we utilize self-attention mechanism (Vaswani et al., 2017) to capture pairwise relations between objectives. Through dynamic computation of relations between objectives, the aligned common parts between them are extracted. Specifically, the softmax normalized dot products between projected objective vectors, $\mathbf{Q}_{r}$ and $\mathbf{K}_{r}$, form the self attention matrix $\mathbf{A}_{r}\in\mathbb{R}^{m\times m}$ between them. This matrix naturally represents the relation strengths between objectives. Then the original score embeddings are integrated into the representations $\mathbf{x}_{r}$ according to the attention matrix.
@@ 4.3. Pre-processing
Before feed the score into relation-aware and relation-agnostic modules. We apply discretization embedding to represent the multi-objective scores to obtain non-linear representation. We use a simple equal distance discretization embedding technique to transform scalar scores to embeddings (Guo et al., 2021).
@@ 5. Experiments
Multi-objective BCE (Cao et al., 2025) 2.2420 3.6465 2.1555 3.6682
Label aggregation (Li et al., 2023) 2.4040 3.8311 2.0765 3.6883
Reinforce learning (Pei et al., 2019) 2.2488 3.9074 2.1930 3.7467
@@ 5.1. Implementation Details
Before training ensemble model, we firstly prepare the input multi-objective scores via performing inference with a high-capacity ranking model over both datasets. For Kuaishou-Elive, we use the online in-service model trained incrementally by years of logs. For the public TenRec dataset, we train a standard MMOE (Ma et al., 2018a) ranking model. Based on prepared multi-objective scores, all offline models are trained with 500 epochs with SGD optimizer. In terms of online training model with streaming data, it is usually trained for one epoch for the consideration of run-time cost and one-epoch overfiting risk in ranking stage. This scheme results in low sample efficiency. Since the model of ME stage is much more lightweight than MP stage and no sparse ID feature is used (cause of one-epoch overfitting), we train 20 epochs for each batch of streaming samples, to balance run-time cost and sample efficiency. We set learning rate to 1e-4 and batch size = 10240 for all methods.
@@ 5.2. Datasets
Since there are few public live-streaming e-commerce dataset, we build a new dataset called Kuaishou E-Live, which contains both e-commerce feedback (purchase) and engagement oriented feedbacks (long view, like, follow and comment). We built it by sampling millions of exposure logs at Kuaishou e-commerce platform.
Moreover, to verify the universality of over method, we conduct experiments on a public short video recommendation dataset, TecRec-QKVideo (Yuan et al., 2022). TecRec-QKVideo is a collected from QQ-KAN video feeds platforms. It contains over 1M user interaction logs including five type of feedbacks, click, like, share, follow and long view. For the need of personalized ensemble, we utilize several user profile features like age, gender, etc. We train our method and compared methods with the same equal loss weights (all weights set to 1.0) for fair comparison on objective trade-offs.
@@
  • Multi-objective binary cross entropy based (M-BCE) (Cao et al., 2025): This method enables optimization towards multiple objectives by integrate joint learning under multiple independent binary cross entropy losses for different objectives.
  • Reinforce learning (RL) based (Pei et al., 2019) : RL-based method formulate multi-objective ensemble as a markov decision process, and treat input multi-objective scores as the ‘state’ and the multi-object fusion score as the ‘action’. The multi-object AUC over a batch of data serve as ‘reward’.
  • Label aggregation based (Carmel et al., 2020; Li et al., 2023) : Label aggregation based method transforms multi-objective learning into single-objective learning by fusion multiple binary labels to a unified regression label. Then it utilize mean square error loss between the ensemble score and the aggregated label to optimize the ensemble model.
@@ 5.4. Offline Results (RQ1)
The experimental results on two datasets are shown in Tab. 1. On the TenRec-QKVideo dataset, our proposed method outperforms MBCE (Cao et al., 2025) by 27.2 pp (3.919 v.s. 3.647) for 5-objective AUC sum. On the Kuaishou-Elive dataset, our model outperforms MBCE by 8.4 pp (3.7522 v.s. 3.6682) for 5-objective AUC sum. The comparison results to the MBCE model demonstrates the efficacy of our proposed ranking-aligned ensemble model.
The RL-based model is the best performing baseline on both datasets. On the TenRec-QKVideo dataset, our model outperforms RL-based model (Pei et al., 2019) by 1.2 pp (3.919 v.s. 3.907) for 5-objective AUC sum. On the Kuaishou-ELive dataset, our model outperforms RL-based model by 0.55 pp (3.7522 v.s. 3.7467). These results verify our continuous gradient-based method is superior to RL policy gradient based methods. In a nutshell, our model consistently outperforms the best-performing baselines on both two datasets in a large margin in terms of different experimental settings.
@@ 5.5. Ablation study (RQ2)
Other than our proposed differentiable ranking based AUC maximization methed, we compare several popular pairwise AUC surrogate losses i.e., pairwise logistic (Gao and Zhou, 2012), pairwise square (Gao et al., 2013) and the state-of-the-art instance-wise AUC maximization method AUCM (Yuan et al., 2021). Pairwise methods maximize AUC objective with logistic or square function surrogate to approximate pairwise $\mathbb{I}(f(\mathbf{x}_{+})\geq f(\mathbf{x}_{-}))$. For pairwise computation nature, these methods perform a $O(n^{2})$ time complexity during training process. Different from them, AUCM transforms pairwise maximization objective to an equivalent instance-wise min-max problem, which reduce the training time complexity to $O(n)$. Notwithstanding these efforts, the inherent inconsistency with the AUC metric persists. We compare the practical training speeds of different losses in Section A of the Appendix.